OpenStudy (anonymous):

find three geometric means between - sqrt 2 and -4sqrt 2

OpenStudy (anonymous):

a. 2, -2sqrt2, 4 b. -2, 2sqrt2,-4 c. 2, 2sqrt2, 4 d. A or C

OpenStudy (anthonyym):

"The Geometric Mean is a special type of average where we multiply the numbers together and then take a square root (for two numbers), cube root (for three numbers) etc." https://www.mathsisfun.com/numbers/geometric-mean.html

OpenStudy (anonymous):

let g1,g2,g3 be G.M's between - sqrt2 and -4 sqrt2 Then \[-\sqrt{2},g1,g2,g3,-4\sqrt{2}are \in G.P.\] let r be common ratio. no.of terms=5 \[a1=-\sqrt{2}\] \[a5=-4\sqrt{2}\]

OpenStudy (anonymous):

\[an=ar ^{n-1}\] \[a5=a1r ^{5-1}=a1r^4\] \[-4 \sqrt2=- \sqrt2 r^4,r^4=\frac{ -4 \sqrt2 }{ -\sqrt2 }=4,(r^2)^2=2^2,r^2=2,r=\sqrt2\]

OpenStudy (anonymous):

\[g1=a1r=-\sqrt2*\sqrt2=-2\] \[g2=g1r=-2*\sqrt2=-2\sqrt2\] \[g3=g2r=-2\sqrt2*\sqrt2=-4\]

OpenStudy (anonymous):

r also can be \[-\sqrt2\] g1=a1r\[=-\sqrt2*-\sqrt2=2\] \[g2=a1r=2*-\sqrt2=-2\sqrt2\] \[g3=g2r=-2\sqrt2*-\sqrt2=4\]